OpenStudy (anonymous):

y varies jointly with p and q. When p = 12 and q = 4, y = 144. Find y when p = 10 and q = 6. Answer y = 180 y = 480 y = 356 y = 288

OpenStudy (anonymous):

y = 180

OpenStudy (anonymous):

\[\frac{12*4}{10*6}=\frac{144}{y}\]

OpenStudy (anonymous):

y=180

OpenStudy (anonymous):

10*6*144=12*4 y = 180

OpenStudy (anonymous):

ty